# Fus-MAE: A cross-attention-based data fusion approach for Masked Autoencoders in remote sensing

### v1

@@ Abstract
This is a preprint for a submission to IGARSS 2024.
Self-supervised frameworks for representation learning have recently stirred up interest among the remote sensing community, given their potential to mitigate the high labeling costs associated with curating large satellite image datasets. In the realm of multimodal data fusion, while the often used contrastive learning methods can help bridging the domain gap between different sensor types, they rely on data augmentations techniques that require expertise and careful design, especially for multispectral remote sensing data. A possible but rather scarcely studied way to circumvent these limitations is to use a masked image modelling based pretraining strategy. In this paper, we introduce Fus-MAE, a self-supervised learning framework based on masked autoencoders that uses cross-attention to perform early and feature-level data fusion between synthetic aperture radar and multispectral optical data - two modalities with a significant domain gap. Our empirical findings demonstrate that Fus-MAE can effectively compete with contrastive learning strategies tailored for SAR-optical data fusion and outperforms other masked-autoencoders frameworks trained on a larger corpus.
@@ 1 Introduction
Multi-modal learning has been attracting increasing attention over the past years, for a vast array of modalities such as RGB-Depth [1] or text-image [2]. Notably, recent research has established theoretical justifications for a performance edge of deep multi-modal learning over unimodal [3]. Within the domain data fusion for remote sensing (RS), two modalities are extensively studied: synthetic aperture radar (SAR) and optical imagery. Indeed, these modalities inherently complement each other: while SAR data offers all-weather and cloud-penetrating capabilities, it suffers from speckle noise, rendering its interpretation challenging. On the other hand, optical data, though subject to weather and seasonal constraints, proposes natural-looking (e.g. RGB) and less noisy images, facilitating interpretation. Hence, their combination proves relevant for tasks such as land cover classification, and opens doors to applications like cloud removal [4] and SAR despeckling [5].
Self-supervised learning (SSL) has stirred substantial interest across diverse machine learning fields, such as natural language processing (NLP) [6, 7] and computer vision [8, 9]. One of its key characteristics is its ability to learn powerful representations without the need for labelled data, which is particularly interesting in the domain of RS, where data annotation can be costly and often requires specific expertise.
The increasing availability of large-scale public SAR-optical datasets such as SeCo [10], BigEarthNet-MM [11], SEN12MS [12] and, more recently, SSL4EO-12 [13] fostered research on SSL approaches for SAR-optical fusion. However, the majority of existing research leans towards contrastive learning [14], which, while effective, presents certain limitations. These include a reliance on data augmentations, which need to be carefully designed to adapt to the specificities of remote sensing images (RSI), as well as the necessity for negative samples, which necessitates a large batch size.
Recent advances in masked image modelling (MIM) [9] set a new state-of-the-art for some visual representation learning tasks. Despite MIM avoiding the aforementionned drawbacks of contrastive learning, to the best of our knowledge, the literature on data fusion for RSI using MIM remains relatively scarce [15, 16]. In this paper, we explore this alternative pretraining approach for SSL data fusion, with our contributions being summarized as following:
@@
    We introduce FusMAE, a self-supervised, MAE-based framework able to perform early-level as well as feature-level data fusion.
@@
    We demonstrate empirically that an early fusion strategy based on cross-attention is the most pre-training SAR-optical data fusion strategy for land cover classification tasks.
@@
    We show that our FusMAE model can compete with some of the most recent contrastive learning approaches tailored for RSI data fusion.
@@ 3 Methodology
Our work is inspired from MultiMAE, a masked autoencoder-based architecture with a proven track record for natural images, able to take different modalities as input [28] with its hybrid-stream architecture. In this section, we describe the Fus-MAE architecture by motivating the need for a multi-task encoder in section 3.1 and a multi-task decoder in section 3.2. Two masking strategies are considered, with detailed provided in section 3.3. The overall architecture is shown in Figure 1.
@@ 3.1 Multi-modal encoder
As in MAE [9], our encoder is a ViT [29] and takes linearly embedded vector representations of patches as input tokens. Let $\textbf{I}_{1}\in\mathbb{R}^{H\times W\times C_{1}}$ and $\textbf{I}_{2}\in\mathbb{R}^{H\times W\times C_{2}}$ be the respective tensor representations of a SAR and an optical satellite image. An intuitive fusion strategy would be to stack SAR and optical RSI data along the channel dimension, and create patches from the obtained tensor. Since this early concatenation technique focuses the entire fusion process onto the single patch projection layer, we hypothesize that it would not be expressive enough to effectively describe cross-modal interactions given the significant domain gap between SAR and optical data. To solve this challenging early fusion task, FusMAE replaces the first encoder block by a ”Cross-attended patch projection” module, which encodes finer-grained multi-modal information into the input tokens. Code and model weights will be made publicly available on GitHub.
@@
The main idea behind the replacement of the first encoder block with this XAttnEncoder block is that, given the large domain gap between the two modalities, feeding both streams of unimodal tokens into the encoder block’s self-attention layer would have resulted in an attention map akin to a block diagonal matrix, with poor cross-modality understanding. On the other hand, the XAttnEncoder block incitates the network to model cross-modal interactions very early, creating tokens with unimodal bias as well as relevant cross-modal information [insert illustrations of attention map either here or in experiments section].
@@ 3.3 Masking strategies
We propose to study two masking strategies: independant masking and consistent masking. Following MAE [9], we apply a 75% masking ratio and sample our patches uniformly across modalities.
@@ 4 EXPERIMENTS AND RESULTS
mAP - 100% labels mAP - 1% labels
S1 S2 S1+S2 S1 S2 S1+S2
ImageNet 70,2 85,5 83,4 55,9 58,5 60,5
Dino-MM [24] 69,7 83,9 84,6 52,7 58,7 60,3
SatViT [15] 75,4 85,6 85,5 52,4 58,5 58,0
Fus-MAE w/ XAD 73,5 87,8 86,9 59,3 60,4 63,1
Fus-MAE w/ XAE&D 75,5 87,9 87,9 57,8 70,0 68,7
Table 1: Mean Average Precision results for the BigEarthNet-MM dataset. S1 = SAR data only. S2 = optical data only. S1+S2 = SAR-optical data fusion. XAD = XAttnDecoder, XAE&D = XAttnEncoder&Decoder.
@@
ImageNet 58,6 90,7 75,2 58,6 60,1
Dino-MM [24] 58,8 91,0 71,4 58,8 59,7
SatViT [15] 59,1 91,8 75,2 59,1 59,9
Fus-MAE w/ XAD 59,8 92,4 75,2 59,8 62,0
Fus-MAE w/ XAE&D 59,6 93,6 71,5 59,6 60,1
@@
Table 2: Classification report for the SEN12MS dataset. S1 = SAR data only. S2 = optical data only. S1+S2 = SAR-optical data fusion. XAD = XAttnDecoder, XAE&D = XAttnEncoder&Decoder
@@
Benchmark setup - To study two different fusion strategies, we pretrained two instances of Fus-MAE: Fus-MAE w/ XAD, which only performs feature-level fusion, and Fus-MAE w/XAE&D, which performs both early and feature-level fusion. We train our models for 100 epochs on the 354,196 images the the BigEarthNet training split, applying the AdamW optimizer with batch size 200. The learning rate is warmed up to $1,5625\times 10^{4}$ for the 10 first epochs, and then decayed with a cosine schedule. We train our models on 2 NVIDIA RTX 3090Ti, for about 60 hours. We set ImageNet initialization as a baseline, and complete our benchmark with two pretrained Transformer-based models that were specifically designed for SAR-optical data fusion: DINO-MM [24] and SatVIT [15], respectively representing recent studies in contrastive learning and masked image modelling.
@@
Multilabel classification - To evaluate the effectiveness of our pretraining strategies, we append a linear classifier head on top of our pretrained encoder. We finetune the model over 10 epochs, using the MultiLabelSoftMargin Loss from PyTorch and the AdamW optimizer, and report the mean average precision (mAP) score over the test split. Table 1 summarizes the outcome of these finetuning experiments, where the classifier is trained on unimodal data (S1 or S2) or multimodal data (S1+S2). Notably, Fus-MAE demonstrates superior performance compared to the other architectures, achieving a substantial increase in mAP of approximately 2 percentage points for S2 and S1+S2, with the Fus-MAE w/XAE&D having a slight edge. It is however noteworthy that, across all architectures, the mAP using only SAR data (S1) is significantly lower than when using S2 or S1+S2, suggesting that all models mainly rely on optical data for their predictions in the S1+S2 scenario. Additionally, to assess the quality of the learned representations under label- and resource-scarce conditions, we perform a linear evaluation on 1% of the training labels. In this scenario, the linear patch projection and the encoder weights are frozen, allowing only the weights of the linear classifier to be learned. We train this classifier for 20 epochs, with a batch size of 128, still utilizing the AdamW optimizer. Results are also reported on Table 1, revealing an even larger performance increase compared to other SSL architectures, highlighting the quality of the learned representations of our models.
@@
Transfer learning - To study the generalization potential, we perform a linear evaluation on another downstream task: unimodal land-cover classification on the SEN12MS dataset [12]. Given the unbalanced nature of this dataset, we computed the precision, the recall and the F1-score using the weighted average method. We train this classifier for 10 epochs, with a batch size of 256, applying the LabelSmoothingCrossEntropy loss from Timm. Results of a benchmark with a similar setup as the previous experiment are reported on Table 2. Our models also outperform other techniques across all tracked metrics, although by a thinner margin.
@@ 5 CONCLUSION
In this paper, we introduce Fus-MAE, a novel SSL framework for SAR-optical data fusion. Based on a MAE architecture, it uses cross-attention between two data streams at different stages to perform early and feature-level data fusion. Our model outperforms recent contrastive learning and MIM-based works on various downstream tasks, demonstrating the effectiveness of using cross-attention to describe cross-modal interactions between modalities with a large domain gap. Further research can be conducted to adapt our cross-attention layers to more than 2 modalities and to balance the prediction reliance of our model more evenly between the modalities.

### v2

@@ Abstract
Selected for IGARSS 2024 Oral. Please cite the version on IEEE Xplore.
Self-supervised frameworks for representation learning have recently stirred up interest among the remote sensing community, given their potential to mitigate the high labeling costs associated with curating large satellite image datasets. In the realm of multimodal data fusion, while contrastive learning methods can help bridge the domain gap between different sensor types, they rely on data augmentation techniques that require expertise and careful design, especially for multispectral remote sensing data. A possible but rather scarcely studied way to circumvent these limitations is to use a masked image modelling based pretraining strategy. In this paper, we introduce Fus-MAE, a self-supervised learning framework based on masked autoencoders that uses cross-attention to perform early and feature-level data fusion between synthetic aperture radar and multispectral optical data - two modalities with a significant domain gap. Our empirical findings demonstrate that Fus-MAE can effectively compete with contrastive learning strategies tailored for SAR-optical data fusion and outperforms other masked-autoencoders frameworks trained on a larger corpus. For replicability, code and weights are provided in this github repository.
@@ 1 Introduction
Multi-modal learning has been attracting increasing attention over the past years, for a vast array of modalities such as RGB-Depth [1] or text-image [2]. In particular, recent research has established theoretical justifications for a performance edge of deep multi-modal learning over unimodal [3]. Within the domain of data fusion for remote sensing (RS), two modalities are extensively studied: synthetic aperture radar (SAR) and optical imagery. Indeed, these modalities inherently complement each other: while SAR data offers all-weather and cloud-penetrating capabilities, it suffers from speckle noise, rendering its interpretation challenging. On the other hand, optical data, though subject to weather and seasonal constraints, proposes natural-looking (e.g. RGB) and less noisy images, facilitating interpretation. Hence, their combination proves relevant for tasks such as land cover classification, and opens doors to applications like cloud removal [4] and SAR despeckling [5].
Self-supervised learning (SSL) has stirred substantial interest in various machine learning fields, such as natural language processing (NLP) [6, 7] and computer vision [8, 9]. One of its key characteristics is its ability to learn powerful representations without the need for labeled data, which is particularly interesting in the domain of RS, where data annotation can be costly and often requires specific expertise.
The increasing availability of large-scale public SAR-optical datasets such as, BigEarthNet-MM [10], SEN12MS [11] and, more recently, SSL4EO-12 [12] fostered research on SSL approaches for SAR-optical fusion. However, the majority of existing research leans towards contrastive learning [13], which, while effective, presents certain limitations. These include a reliance on data augmentations, which need to be carefully designed to adapt to the specificities of remote sensing images (RSI), as well as the necessity for negative samples, which necessitates a large batch size, hence large compute resources.
Recent advances in masked image modelling (MIM) [9] set a new state-of-the-art for some visual representation learning tasks. Despite MIM avoiding the above-mentioned drawbacks of contrastive learning, to the best of our knowledge, the literature on data fusion for RSI using MIM remains relatively scarce [14, 15]. In this paper, we explore this alternative pretraining approach for SSL data fusion, with our contributions being summarized as following:
@@
    We introduce Fus-MAE, a self-supervised, MAE-based framework able to perform early-level as well as feature-level data fusion.
@@
    We demonstrate empirically that an early-fusion approach leveraging cross-attention is the best pre-training strategy for transformers to perform SAR-optical data fusion tasks.
@@
    We show that our Fus-MAE model can compete with some of the most recent contrastive learning approaches tailored for RSI data fusion.
@@ 3 Methodology
Our work is inspired by MultiMAE, a masked autoencoder-based architecture with a proven track record for natural images, capable of taking different modalities as input [27] with its hybrid-stream architecture. In this section, we describe the Fus-MAE architecture by motivating the need for a multi-task encoder in section 3.1 and a multi-task decoder in section 3.2. Two masking strategies are considered, with detailed provided in section 3.3. The overall architecture is shown in Figure 1.
@@ 3.1 Multi-modal encoder
As in MAE [9], our encoder is a ViT [28] and takes linearly embedded vector representations of patches as input tokens. Let $\textbf{I}_{1}\in\mathbb{R}^{H\times W\times C_{1}}$ and $\textbf{I}_{2}\in\mathbb{R}^{H\times W\times C_{2}}$ be the respective tensor representations of a SAR and an optical satellite image. An intuitive fusion strategy would be to stack SAR and optical RSI data along the channel dimension, and create patches from the obtained tensor. Since this early concatenation technique focuses the entire fusion process onto the single patch projection layer, we hypothesize that it would not be expressive enough to effectively describe cross-modal interactions given the significant domain gap between SAR and optical data. To solve this challenging early fusion task, Fus-MAE replaces the first encoder block by a ”Cross-attended patch projection” module, which encodes finer-grained multi-modal information into the input tokens.
@@
The main idea behind the replacement of the first encoder block with this XAttnEncoder block is that, given the large domain gap between the two modalities, feeding both streams of unimodal tokens into the encoder block’s self-attention layer would have resulted in an attention map akin to a block diagonal matrix, with poor cross-modality understanding. On the other hand, the XAttnEncoder block incitates the network to model cross-modal interactions very early, creating tokens with unimodal bias as well as relevant cross-modal information.
@@ 3.3 Masking strategies
We propose to study two masking strategies: independent masking and consistent masking. Following MAE [9], we apply a 75% masking ratio and sample our patches uniformly across modalities.
@@ 4 EXPERIMENTS AND RESULTS
S1 S2 S1+S2
ImageNet 70.2 85.5 83.4
Dino-MM 69.7 83.9 84.6
SatViT 75.4 85.6 85.5
Fus-MAE XAD 75.1 86.9 87.2
Fus-MAE XAE 75.9 87.6 88.1
S1 S2 S1+S2
ImageNet 55.9 58.5 60.5
Dino-MM 52.7 58.7 60.3
SatViT 52.4 58.5 58.0
Fus-MAE XAD 64.8 71.8 72.2
Fus-MAE XAE 57.5 68.0 70.0
Table 1: Mean Average Precision results for the BigEarthNet-MM dataset. Left = finetuning with 100% of labels. Right = linear evaluation with 1% labels. S1 = SAR data only. S2 = optical data only. S1+S2= SAR-optical data fusion. XAD = XAttnDecoder, XAE = XAttnEncoder
@@
ImageNet 58.6 90.7 75.2 58.6 60.1
Dino-MM 58.8 91.0 71.4 58.8 60.3
SatViT 59.1 91.8 75.2 59.1 59.9
Fus-MAE XAD 58.8 94.0 73.9 58.7 60.5
Fus-MAE XAE 60.6 94.3 71.5 60.6 61.0
@@
Table 2: Classification report for the linear evaluation experiment on the SEN12MS dataset. XAD = XAttnDecoder, XAE = XAttnEncoder.
@@
Benchmark setup - To study two different fusion strategies, we pretrained two instances of Fus-MAE: Fus-MAE XAE, which performs early-level fusion during encoding, and Fus-MAE XAD, which performes feature-level fusion during decoding. We train our models for 100 epochs on the 354,196 images the the BigEarthNet training split, applying the AdamW optimizer with batch size 200, with a learning rate of $1,5625\times 10^{4}$. We train our models on 2 NVIDIA RTX 3090Ti, for about 60 hours. We set ImageNet initialization as the baseline, and complete our benchmark with two pretrained Transformer-based models that were specifically designed for SAR-optical data fusion: DINO-MM [23] and SatVIT [14], respectively representing recent studies in contrastive learning and masked image modelling.
@@
Multilabel classification - To evaluate the effectiveness of our pretraining strategies, we append a linear classifier head on top of our pretrained encoder. We finetune the model over 10 epochs, using a multi-label soft margin loss and the AdamW optimizer, and report the mean average precision (mAP) score over the test split. Table 1 summarizes the outcome of these finetuning experiments, where the classifier is trained on unimodal data (S1 or S2) or multimodal data (S1+S2). On this task, the early fusion approach shows higher performance, showing the potency of our cross-attended patch projection module. It is noteworthy that, across all architectures, the mAP using only SAR data (S1) is significantly lower than when using S2 or S1+S2, suggesting that all models mainly rely on optical data for their predictions in the S1+S2 scenario. Additionally, to assess the quality of the learned representations under label- and resource-scarce conditions, we perform a linear evaluation on 1% of the training labels. In this scenario, the linear patch projection and the encoder weights are frozen, allowing only the weights of the linear classifier to be learned. We train this classifier for 20 epochs, with a batch size of 128, still utilizing the AdamW optimizer. Results are also reported on Table 1, revealing an even larger performance increase compared to other SSL architectures, highlighting the quality of the learned representations of our models.
@@
Transfer learning - To study the generalization potential, we perform a linear evaluation on another downstream task: unimodal land-cover classification on the SEN12MS dataset [11]. We train this classifier for 10 epochs. Given the unbalanced nature of this dataset, we computed the classification metrics using the weighted average method, and applied a label smoothing cross entropy loss. Results of a benchmark with a similar setup as the previous experiment are reported on Table 2. Our models also outperform other techniques across all tracked metrics, although by a thinner margin.
@@ 5 CONCLUSION
In this paper, we introduce Fus-MAE, a novel SSL framework for SAR-optical data fusion. Based on the MAE architecture, it uses cross-attention between two data streams at different stages to perform early and feature-level data fusion. Our model outperforms recent contrastive learning and MIM-based works on various downstream tasks, demonstrating the effectiveness of using cross-attention to describe cross-modal interactions between modalities with a large domain gap. Further research can be conducted to adapt our cross-attention layers to more than 2 modalities and to balance the prediction reliance of our model more evenly between the modalities.